# LLMs are Biased Teachers: Evaluating LLM Bias in Personalized Education

### v1

@@ 5.1 RQ1: Bias in Educational Text Selection
Table 2: Selected examples of bias in generated explanations from GPT 4o. Row 1: The model assumes that the Hispanic student has personal connections to border security, which it does not do for any other race/ethnicity. Row 2: The model presents highly religious content that may distract from the subject. Row 3: The model presents an example of fruit for the low-income student while assuming international business knowledge from the high-income student. It is important to note these overt biases could be even more harmful if the model assumes demographic information incorrectly.
Bias Magnitude.
These results indicate substantial biases across all subgroups, with Disability Status and Income [^7] Notably, the Income subgroup closely matches the bias patterns of the Reference classes in nearly all experiments. receiving the highest bias levels out of the tested subgroups. We also find that the bias magnitude for each model varies based on subgroup and metric, and there is no definitive “most biased” of our tested LLMs. As for datasets, News In Levels consistently produces the highest bias scores.
Refusals.
We encounter two types of refusals in the model outputs during ranking. First are partial refusals, where the LLM states that it would be wrong to base the explanation on the student class but still provides an output. We find that these do not significantly influence the overall bias results. Second are full refusals, where the model does not select any explanation. We only observe full refusals with Llama 3.1 405B and note that the refusal pattern itself is biased (>90% refused for “black” and “native american” and <20% for all others). We filter full refusals from our data, which is reflected with larger error bars.
5.2 RQ2: Student vs Teacher Personas
In this section, we investigate if there are differences in ranking bias between our setting (“teacher persona”) and the traditional persona setting (“student persona”). We find that biases perpetuate with similar patterns for both persona settings.
Using the prompt described in Section 4.4, we run additional ranking experiments. We plot the teacher and student persona data for selected subgroups in Figure 4 and complete results in Appendix D.5. We observe that the biases in the student persona generally mirror trends seen in the teacher persona, especially in the Race/Ethnicity and Gender/Sexuality categories. In other models, we see some variation in the Religion, Income, and National Origin categories, but in GPT 4o, trends align consistently across all categories. This demonstrates the LLM does not differentiate the student and the teacher personas, relying solely on demographic characteristics to determine the appropriate level of explanation.
(a) Bias in the Math-50 ranking task.
(b) Bias in the 7 topics ranking task.
Figure 5: The normalized scores (3) and 95% bootstrapping CI of GPT 4o for MATH and topic modeling, selected subgroups of interest. In (b), we observe that the subject of the articles does not seem to have much of an effect on the models’ choices. Most variation falls within the error bars, and trends follow the overall News In Levels experiments. In (a), we observe similar bias patterns across most categories despite the lack of linguistic features. In both figures, we observe similar instances of reverse bias (e.g. Disability)
@@ 5.3 RQ3: Bias in the Generative Task
Compared with the ranking task, we observe bias scores for generation at a level between the WIRED and machine-generated tasks, as shown in Figure 3. Additionally, Alignment with the reference classes is more substantial (Figure 2(b)). In fact, for the classes from the WIRED videos, we see strong alignment with the grade levels (“child” $\approx 7$, “teen” $\approx 10$, “college student” $\approx 13$, “graduate” $\approx 17$, “expert” $\approx 20$).
@@
Overt Bias in Generations.
In addition to bias in the linguistic complexity of explanations, we also observe explicit stereotypes in the generations. Selected examples of these are available in Table 2. Additionally, we note that all models will often respond to the Hispanic student in Spanish despite providing English responses for all other classes and the prompt being in English.
@@ 5.4 RQ4: Topic-Specific Biases
We explore the effect of individual topics on model bias through topic modeling and do not find evidence of topic-specific bias patterns. We manually cluster News In Levels articles into 7 topics and select the corresponding outputs from each model for each article in a topic. This forms grouped subsets of the results for each topic, which we use to visualize the ranking trends. The results for selected subgroups from the 7 topic experiments on GPT 4o are presented in Figure 5(b), while other models can be found in Appendix D.6.
We also try BertTopic Grootendorst (2022) for clustering in 4 to 70 topics, which still does not provide evidence of a topic-wise difference (Appendix D.6). Experiments with fewer topics show less variation between topics, while more topics show higher variation but also larger error margins.
@@
In this setting, we explore if bias patterns are prevalent in the mathematical setting, where linguistic complexity does not correlate with difficulty. We provide each model problem-solution pairs from the MATH-50 dataset, as described in Section 4.1 with the same prompt as Section 5.1. Surprisingly, similar bias patterns propagate, including instances of reverse bias discussed in Section 5.1. The results can be found in Figure 5(a) and in Appendix D.7.
@@ 6 Discussion and Conclusion
This study evaluates the biases in LLMs when used as a “teacher persona” in personalized education. We find significant biases across demographic groups, such as race, gender, disability, and income. These biases persist in ranking and generation, for multiple persona types, and across diverse topics. The impact of these biases is clear—LLMs risk perpetuating harmful stereotypes and reinforcing inequality in educational outcomes. As LLMs are increasingly integrated into personalized learning environments, addressing these biases becomes critical to avoid exacerbating existing disparities.
To ensure fair and inclusive AI in education, future work must focus on mitigating these biases by improving LLM training and evaluation processes. This includes developing methods that adapt content based on students’ learning progress and ignore demographic assumptions. Research should also explore how these biases affect learning outcomes over time, particularly for marginalized groups. By addressing these challenges, we can better align LLMs with equitable and effective education.
@@
While our study explores the biases of LLMs in the context of personalized teaching, it is important to acknowledge that we are not directly evaluating LLMs designed explicitly for personalized education. Instead, we use state-of-the-art Foundation LLMs as proxies by assigning teacher and student personas with explicit protected attributes. This approach has been discussed in-depth in the literature Bommasani et al. (2022). In real-world scenarios, however, protected attributes may not always be explicitly provided. Nonetheless, as discussed in the main text, LLMs can infer such attributes through indirect means, such as names, school locations, prior interactions, or even past student-written work, which may introduce unintentional bias.
@@
We also emphasize that we do not intend to dissuade the use of LLMs for education, even in the personalized setting, and instead call for further study into the possible harms and measures to prevent them. When referencing “common” or “typical stereotypes” we follow those available in Nadeem et al. (2021) but recognize that these vary between cultures and situations.
@@ B.5 Generated Dataset using WIRED subjects
Appendix C LLMs and Experiments
@@
Model Generative Persona Topics
@@
Table 3: Models, specific versions, and their generative, student persona, and topic modeling experiments
@@
Appendix D Additional Plots
D.1 Ranking/WIRED Experiments
@@
D.2 Ranking/NewsinLevels Experiments
@@
D.3 Ranking/Generated Diverse Experiments
@@
D.4 Ranking/Generated WIRED Experiments
@@
D.5 Ranking/Student Persona Experiments
Figure 32: Persona Bias plots for NewsInLevels dataset on Gemini-1.5.
Figure 33: Persona Bias plots for News In Levels dataset on GPT 4 Turbo.
Figure 34: Persona Bias plots for News In Levels dataset on GPT 4o.
D.6 Ranking/Topic Modeling Experiments
@@
D.7 Ranking/MATH-50 Experiments
@@
D.8 Generative/WIRED Subjects

### v2

@@ 5.2 RQ2: Student vs Teacher Roles
Table 2: Selected examples of generated explanations from GPT 4o. We see explicit stereotyping and overpersonalization, which both may distract from the subject and harm student learning. Row 1: The model assumes that the Hispanic student has personal connections to border security, which it does not do for any other race/ethnicity Flores and Schachter (2019). Row 2: The model presents highly religious content that may distract from the subject. Row 3: The model presents an example of fruit for the low-income student while assuming international business knowledge from the high-income student. It is important to note these overt biases could be even more harmful if the model assumes demographic information incorrectly.
@@ 5.3 RQ3: Bias in the Generative Task
Compared with the ranking task, we observe bias scores for generation at a level between the WIRED and machine-generated tasks, as shown in Figure 3. Additionally, Alignment with the reference characteristics is more substantial (Figure 2(b)). In fact, for the characteristics from the WIRED videos, we see strong alignment with the grade levels (“child” $\approx 7$, “teen” $\approx 10$, “college student” $\approx 13$, “graduate” $\approx 17$, “expert” $\approx 20$).
@@
Stereotypes in Generations.
In addition to bias in the linguistic complexity of explanations, we also observe explicit stereotypes in the generations. Selected examples of these are available in Table 2. Additionally, we note that all models will often respond to the Hispanic student in Spanish despite providing English responses for all other characteristics and the prompt being in English. On top of stereotypes, we observe instances of extreme overpersonalization to demographic traits, which are irrelevant to the subject and may distract from student learning.
@@ 5.4 RQ4: Topic-Specific Biases
We explore the effect of individual topics on model bias through topic modeling and do not find evidence of topic-specific bias patterns. We manually cluster News In Levels articles into 7 topics and select the corresponding outputs from each model for each article in a topic. This forms grouped subsets of the results for each topic, which we use to visualize the ranking trends. The results for selected subgroups from the 7 topic experiments on GPT 4o are presented in Figure 5(b), while other models can be found in Appendix E.6.
We also try BertTopic Grootendorst (2022) for clustering in 4 to 70 topics, which still does not provide evidence of a topic-wise difference (Appendix E.6). Experiments with fewer topics show less variation between topics, while more topics show higher variation but also larger error margins.
@@
In this setting, we explore if bias patterns are prevalent in the mathematical setting, where linguistic complexity does not correlate with difficulty. We provide each model problem-solution pairs from the MATH-50 dataset, as described in Section 4.1 with the same prompt as Section 5.1. Surprisingly, similar bias patterns propagate, including instances of reverse bias discussed in Section 5.1. The results can be found in Figure 5(a) and in Appendix E.7.
@@ 6 Discussion and Conclusion
This study evaluates the biases in LLMs when used as a “teacher” in personalized education. We find significant biases across demographic groups, such as race, gender, disability, and income. These biases persist in ranking and generation, for multiple roles, and across diverse topics. The impact of these biases is clear—LLMs risk perpetuating harmful stereotypes and reinforcing inequality in educational outcomes. As LLMs are increasingly integrated into personalized learning environments, addressing these biases becomes critical to avoid exacerbating existing disparities.
We note that this issue is increasingly urgent as LLM-based educational tools are rapidly being deployed across K-12 and college-level education through platforms like Cognii, Duolingo, and Khan Academy. These tools may disproportionately benefit majority (“average”) students while potentially harming underrepresented students through biased treatment. While many education software may not have access to students’ protected attributes directly, it is well-documented that LLMs can easily infer protected attributes from students’ names Haim et al. (2024); You et al. (2024); Nghiem et al. (2024), school zipcodes Jeong et al. (2022), language Levy et al. (2023), and other cues Kantharuban et al. (2024). Hence, the biased behavior we document in this paper can easily be replicated in real-world settings.
Our ranking experiment can map to “grounded” retrieval-based systems that have access to human-written materials (ex. Khan Academy). We observed substantial bias in this setting, implying that some demographic groups can get inaccurate assignment of explanation difficulty level. A body of psychology literature emphasizes the “desirable difficulty” level for optimizing learning Vygotskij and Cole (1981); Csikszentmihalyi (2009). Long-term effects on students due to biased LLM teachers is an important future research area.
In the generative setting, we observe stereotypes embedded in the generated explanation, which goes beyond inaccurate difficulty level assignment. In the literature on generative personalization, there is an active open discussion about the appropriate level of personalization Gautam et al. (2024); Kantharuban et al. (2024). Specifically, what degree of stereotyping crosses the line into harmful territory and what can be thought of as accommodation? In the education setting, having a relatable teacher correlates with students’ success, especially students from marginalized groups Keane et al. (2023); Thiem and Dasgupta (2022). We believe that substantially more future work should be done to understand a clear line between stereotyping LLM teachers and relatable LLM teachers.
@@
While our study explores the biases of LLMs in the context of personalized teaching, it is important to acknowledge that we are not directly evaluating LLMs designed explicitly for personalized education. Instead, we use state-of-the-art Foundation LLMs as proxies by assigning teacher and student roles with explicit protected attributes. This approach has been discussed in-depth in the literature Bommasani et al. (2022). In real-world scenarios, however, protected attributes may not always be explicitly provided. Nonetheless, as discussed in the main text, LLMs can infer such attributes through indirect means, such as names, school locations, prior interactions, or even past student-written work, which may introduce unintentional bias.
@@
We also emphasize that we do not intend to dissuade the use of LLMs for education, even in the personalized setting, and instead call for further study into the possible harms and measures to prevent them. When referencing “common” or “typical stereotypes” we follow those referenced in Section 2 but recognize that these vary between cultures and situations.
Acknowledgments
This work was supported in part by the National Science Foundation under Grant No. 2341055.
@@ B.5 Generated Dataset using WIRED subjects
Appendix C Details on Readability Scores
When calculating the linguistic complexity using the three grade level metrics (Equation (2), we find that all three are strongly linearly correlated for our data (Pearson $R>0.96$ for all metrics and TGL) and map to the same scale (US Grade Levels). Thus, because of these similarities, we opted to take the mean rather than selecting a single metric (as the field lacks consensus) or plotting all metrics individually (which reduced the interpretability of plots). Matsuura and Jaeah (2024) provides a more detailed comparison of popular traditional readability metrics in the early education context, also finding agreement between the three we use.
The three individual metrics are calculated as following:
The Flesch-Kincaid Grade Level Kincaid et al. (1975) formula:
FKGL=0.39(total wordstotal sentences)+11.8(total syllablestotal $\text{FKGL}=0.39\left(\frac{\text{total words}}{\text{total sentences}}\right)+11.8\left(\frac{\text{total syllables}}{\text{total words}}\right)-15.59$
The Gunning Fog Index Gunning (1952) formula (“complex words” are those with three or more syllables):
$\text{Fog}=0.4\left[\left(\frac{\text{words}}{\text{sentences}}\right)+100\left(\frac{\text{complex words}}{\text{words}}\right)\right]$
The Coleman-Liau Index Coleman and Liau (1975) formula:
$\text{CLI}=0.0588\left(\frac{100*characters}{words}\right)-0.296\left(\frac{100*sentences}{words}\right)-15.8$
Appendix D LLMs and Experiments
@@
Model Generative Student Role Topics
@@
Table 3: Models, specific versions, and their generative, student role, and topic modeling experiments
@@
Appendix E Additional Plots
E.1 Ranking/WIRED Experiments
@@
E.2 Ranking/NewsinLevels Experiments
@@
E.3 Ranking/Generated Diverse Experiments
@@
E.4 Ranking/Generated WIRED Experiments
@@
E.5 Ranking/Student Role Experiments
Figure 32: Student Role plots for NewsInLevels dataset on Gemini-1.5.
Figure 33: Student Role plots for News In Levels dataset on GPT 4 Turbo.
Figure 34: Student Role plots for News In Levels dataset on GPT 4o.
E.6 Ranking/Topic Modeling Experiments
@@
E.7 Ranking/MATH-50 Experiments
@@
E.8 Generative/WIRED Subjects